# MERLIN: Multi-stagE query performance prediction for dynamic paRallel oLap pIpeliNe

### v1

@@ 5.2. Cost Calibration
Figure 4. Illustration of ClickHouse’s physical operators’ relative resource cost. The red operators are marked as existing corresponding resource competition. All ‘-ing’ and ‘Transform’ post-fix are omitted for clarity.
Resource Competition Attention. Although the resource competition matrices are summarized based on both operators’ code analysis results and relative cost evaluation, resource contention between operators with resource competition relations may not occur during the pipeline execution process. Even if it does occur, the intensity of resource contention may vary depending on system load, pipeline structure, data characteristics, operator call parameters, etc. The existing work (Zhou et al., 2020) uses estimated operators’ latency to build a timeline of operators’ calling and estimate the competition intensity by the operators’ overlapped ratio. However, such a solution is vulnerable and the error would accumulate with the growth of the operator’s number since each operator’s start and end time depend on every operator’s latency before it. For high-performance OLAP DBMSs like ClickHouse and SparkSQL, the pipeline is usually longer and involves parallel branches. Besides, they execute operators immediately once the operator is ready to be executed. So when we dive to the level of each step, the execution time and order of operators are more random, so it is impossible to accurately predict the timeline and overlap the time between operators like in the existing work (Zhou et al., 2020). Instead, we introduce an attention mechanism to self-adjust operator competition intensity based on the given data-flow tree and resource competition DAG.
The core idea of attention is to calculate the attention score that represents the correlation between every item in features $Q,K$ and apply that score to the feature $V$ (Vaswani et al., 2017).
(2) $Attention(Q,K,V)=softmax(\frac{QK^{\mathsf{T}}}{\sqrt{d_{k}}})V$
In this case, MERLIN needs to calculate the operator’s resource competition intensity based on the given data-flow tree $\mathcal{M}_{F}$. Thus, the attention score should be calculated from the resource competition matrix $\mathcal{M}^{resource}_{F}$ and the data-flow tree $\mathcal{M}_{F}$, which can be formalized as:
(3) $\mathcal{M^{\prime}}^{resource}_{F}=softmax(\frac{\mathcal{M}^{resource}_{F}\mathcal{M}_{F}^{T}}{\sqrt{N}})\mathcal{M}^{resource}_{F}$
@@
After constructing the graph, MERLIN uses a GAT model (Veličković et al., 2017) to capture the resource competition and calibrate the operators’ cost vectors. Each layer of the GAT model can be formalized with the following equation:
(4) $x^{\prime}_{i}$ $=\sum_{j\in\mathcal{N}(i)\cup\{i\}}{\alpha_{i,j}\Theta_{t}x_{j}}$
(5) $\alpha_{i,j}$ $=\frac{exp(\alpha^{\mathsf{T}}LeakyReLU(\Theta_{s}x_{i}+\Theta_{t}x_{j}+\Theta_{e}e_{i,j}))}{\sum_{k\in\mathcal{N}(i)\cup\{i\}}{exp(\alpha^{\mathsf{T}}LeakyReLU(\Theta_{s}x_{i}+\Theta_{t}x_{k}+\Theta_{e}e_{i,k}))}}$
, where $e_{i,j}$ is the elements in $\mathcal{M}_{pipeline}$ that represents the connectivity and the resource competition intensity.
After being processed using all the methods above, MERLIN now gets the calibrated resource cost vector of each operator in the pipeline. MERLIN then converts the calibrated resource vectors back to the pipeline tree structure and summarizes those vectors from the bottom to the top by a TCN to predict the query performance. Note that all learnable weights in this section are trained by QPP tasks together with TCN introduced in the next subsection.
5.3. Query Performance Prediction
In this section, we introduce how MERLIN predicts the query performance based on calibrated cost vectors and how to train the calibrator and the predictor together.
In subsection 5.2, we express the pipeline as the data-flow tree and each operator’s predicted cost vector is calibrated according to the DBMS’s operator parallelization mechanism. Considering the high-performance OLAP DBMSs’ execution mechanisms, simply adding each operator’s cost vector as the final predicted performance is too rough. Instead, MERLIN combines those calibrated cost vectors according to the data-flow tree structure to construct a vector tree. We use a tree convolutional network (TCN) (Mou et al., 2016) to summarize the cost from bottom to top. The TCN has two main advantages. Firstly, it is powerful to capture tree shape query plan’s feature with a low inference cost, which is the reason why it has been widely used in many AI4DB methods (Marcus et al., 2019; Lan et al., 2021). Secondly, the paper of TCN proposed a continuous binary tree mode, which native supports trees with multiple children nodes. Since the data-flow tree is constructed from execution paths of a DAG pipeline, many nodes have more than two children, and TCN can handle such a scenario well.
The TCN has a triangle shape convolution filter that is made of three weight matrices $W^{t}_{conv},W^{l}_{conv},W^{r}_{conv}$. For node $x_{i}$ in the window of a filter, its weight matrix is denoted as $W_{conv,i}$, which is a linear combination of those three weights above with corresponding coefficients $\eta_{i}^{t},\eta_{i}^{r}$, and $\eta_{i}^{l}$. Each layer’s output of TCN is calculated by the following formulas (Mou et al., 2016):
(6) $\eta_{i}^{t}=\frac{d_{i}-1}{d-1}\quad\eta_{i}^{r}$ $=(1-\eta_{i}^{t})\frac{p_{i}-1}{n-1},\quad\eta_{i}^{l}=(1-\eta_{i}^{t})(1-\eta_{i}^{r})$
(7) $W_{conv,i}$ $=\eta_{i}^{t}W^{t}_{conv}+\eta_{i}^{l}W^{l}_{conv}+\eta_{i}^{r}W^{r}_{conv}$
(8) $y$ $=tanh(\sum^{n}_{i=1}{W_{conv,i}\cdot x_{i}+b_{conv}})$
@@ 5.3. Query Performance Prediction
MERLIN’s GAT and TCN networks are trained together in a query-driven way since constructing the calibrated cost vector to the cost vector tree is differentiable. They are trained with the MSE loss function together with normalized ground-truth query latency as the label. To make the model converge more stable, all neural networks are connected by residual connection path (He et al., 2016).
6. Experiments
@@
In this section, we evaluate MERLIN’s performance and compare it with other baseline methods. Since ClickHouse supports all the mechanisms as shown in Table 1, we consider it one of the most representative high-performance OLAP DBMSs. Thus, we implement MERLIN on ClickHouse for evaluation.
6.1. Experimental Settings
@@
Baselines. Since previous works already applied full comparison with conventional methods such as SVM and RBF (Marcus and Papaemmanouil, 2019; Akdere et al., 2012b; Li et al., 2012b) and proved learned QPP approaches out-performed conventional approaches on average accuracy, we will focus on the deep learning-based methods proposed in recent years. All baselines compared are listed by following:
  1. (1)
@@
    QPPNet (Marcus and Papaemmanouil, 2019): The QPPNet is an operator-level method that builds a cost model with a unified input and output interface for each operator and combines them according to the structure of the query plan tree to predict the latency. This method is considered to be a flexible model that can adapt to all different plan templates.
  2. (2)
@@
    QueryFormer (Zhao et al., 2022): The QueryFormer is a variant of the Transformer, which provides an attention mechanism on tree-shaped data and can capture a high-level expression of query plans to support downstream tasks such as QPP.
  3. (3)
@@
    RAAL (Li et al., 2022): The RAAL uses an attention-based resource-aware network to capture the resource allocation in SparkSQL and predict the query latency. The application scenario is quite similar to ClickHouse’s parallel pipeline so we include it as a competitive baseline. We implement it by following the formula and instructions in its paper.
  4. (4)
@@
    GCN  (Kipf and Welling, 2016): The GCN is a well-known network that is proposed to perform convolution on graph data. Since both DAG and tree-shaped pipelines can be regarded as a graph, we use GCN to perform end-to-end QPP as a baseline.
  5. (5)
@@
    TCN  (Mou et al., 2016): The TCN is a common and widely used network in the domain related to query execution and query optimization (Marcus et al., 2019), which introduces convolution operation to the tree-shaped data. Thus, we included it as a baseline.
@@
We used their open-source code and modified it to adapt the data collected by MERLIN’s Runtime Tracker from ClickHouse due to ClickHouse’s lack of a tool to export a detailed enough physical pipeline for baselines such as QPPNet and QueryFormer. In fact, MERLIN’s Runtime Tracker provides even more comprehensive features than their original implementation. For RAAL, since it is an approach designed for SparkSQL that considers the resource competition, which is also included in MERLIN’s scenario, it can share the same training and testing data with MERLIN (only in a different format). Thus, the comparison is carefully designed and implemented to ensure fairness.
@@
In addition to those baselines above, we also performed ablation experiments on MERLIN to evaluate each component’s performance with the following setups:
  1. (1)
@@
    MERLIN w/o ResAttn: This combination only removes the resource attention mechanism of MERLIN and uses the expanded meta-competition matrix only as GAT’s input.
  2. (2)
@@
    MERLIN w/o OCP: This combination removes the OCP module only and keeps the other modules to evaluate the impact of the OCP module on the QPP performance.
@@
Datasets. Followed by QPPNet (Marcus and Papaemmanouil, 2019) and MB2 (Ma et al., 2021), we compare the performance on TPC-H and TPC-DS benchmarks with 1GB scale factor. These are two standard benchmarks designed specifically for evaluating OLAP database performance and be widely used in the industry.
@@
Workloads. After removing unsupported query templates, 17 templates for TPC-H and 58 templates for TPC-DS are left for training and testing, the number of generated queries of each workload is about 10,000 (slightly fluctuating based on the number of query templates because we generate the same number of queries for each template). Among those queries, we evenly draw all the queries of 4 templates (templates 3, 7, 11, 6) from TPC-H (23.5%) and 9 templates (templates 3, 9, 15, 21, 27, 33, 39, 45, 51, 57) from TPC-DS (15.52%) by the index as test workloads that will not appear in the training workloads. Note that the testing workloads only use the data collected by the probe execution mode as input.
@@
Metric. We use the Q-Error as the metrics to measure the accuracy of QPP, which is widely used in many previous QPP and cardinality estimation methods (Marcus and Papaemmanouil, 2019; Wu et al., 2013b; Yang et al., 2019). The definition of Q-Error is listed as follows:
(9) Q-Error $=\frac{\max{(prediction,actual)}}{\min{(prediction,actual)}}$
@@
Model Architecture. Considering the inference speed, MERLIN uses relatively shallow networks, such as a 2-layer GAT and a 3-layer TCN and all OCPs are MLP with 3 hidden layers.
@@
Hardware. All training and testing workloads are generated on AMD R9-7945HX, 32GB memory, and RTX4060 with a controllable environment that has no interference from other processes. We use a public server with 6 RTX A6000 (48GB memory) GPUs, 2 Intel(R) Xeon(R) Silver 4210R CPUs, and 504GB RAM to parallelly perform all evaluations on MERLIN, baselines, and ablation experiments.
6.2. QPP Accuracy
Table 2. Accuracy comparison of baselines, MERLIN, and ablation studies.

### v2

@@ IV-B Cost Calibration
Figure 2: Illustration of ClickHouse’s physical operators’ relative resource cost. The red operators are marked as existing corresponding resource competition.
Resource Competition Attention. Although the resource competition matrices are summarized based on both operators’ code analysis results and relative cost evaluation, resource contention between operators with resource competition relations may not occur during the pipeline execution process. Even if it does occur, the intensity of resource contention may vary depending on system load, pipeline structure, data characteristics, operator call parameters, etc. The existing work [12] uses estimated operators’ latency to build a timeline of operators’ calling and estimate the competition intensity by the operators’ overlapped ratio. However, such a solution is vulnerable, and the error would accumulate with the growth of the operator’s number since each operator’s start and end time depend on every operator’s latency before it. For high-performance OLAP DBMSs, the pipeline is usually longer and involves parallel branches. Besides, they execute operators immediately once the operator is ready to be executed. So when we dive to the level of each step, the execution time and order of operators are more random, so it is impossible to accurately predict the timeline and overlap the time between operators like in the existing work [12]. Instead, we introduce an attention mechanism to self-adjust operator competition intensity based on the given data-flow tree and resource competition DAG.
The core idea of attention is to calculate the attention score that represents the correlation between every item in features $Q,K$ and apply that score to the feature $V$ [29].
$Attention(Q,K,V)=softmax(\frac{QK^{\mathsf{T}}}{\sqrt{d_{k}}})V$ (2)
In this case, CONCERTO needs to calculate the operator’s resource competition intensity based on the given data-flow tree $\mathcal{M}_{F}$. Thus, the attention score should be calculated from the resource competition matrix $\mathcal{M}^{resource}_{F}$ and the data-flow tree $\mathcal{M}_{F}$, which can be formalized as:
$\mathcal{M^{\prime}}^{resource}_{F}=softmax(\frac{\mathcal{M}^{resource}_{F}\mathcal{M}_{F}^{T}}{\sqrt{N}})\mathcal{M}^{resource}_{F}$ (3)
@@
After constructing the graph, CONCERTO uses a GAT model [21] to capture the resource competition and calibrate the operators’ cost vectors.
After being processed using all the methods above, CONCERTO now gets the calibrated resource cost vector of each operator in the pipeline. CONCERTO then converts the calibrated resource vectors back to the pipeline tree structure and summarizes those vectors from the bottom to the top by a TCN to predict the query performance. Note that all learnable weights in this section are trained by QPP tasks together with TCN introduced in the next subsection.
IV-C Query Performance Prediction
In this section, we introduce how CONCERTO predicts the query performance based on calibrated cost vectors and how to train the calibrator and the predictor together.
In subsection IV-B, we express the pipeline as a data-flow tree, and each operator’s predicted cost vector is calibrated according to the DBMS’s operator parallelization mechanism. Considering the high-performance OLAP DBMSs’ execution mechanisms, simply adding each operator’s cost vector as the final predicted performance is too rough. Instead, CONCERTO combines those calibrated cost vectors according to the data-flow tree structure to construct a vector tree. We use a tree convolutional network (TCN) [30] to summarize the cost from bottom to top. The TCN has two main advantages. Firstly, it is powerful to capture tree shape query plan’s feature with a low inference cost, which is the reason why it has been widely used in many AI4DB methods [4, 28]. Secondly, the paper of TCN proposed a continuous binary tree mode, which natively supports trees with multiple child nodes. Since the data-flow tree is constructed from execution paths of a DAG pipeline, many nodes have more than two children, and TCN can handle such a scenario well.
The TCN has a triangle-shaped convolution filter that is made of three weight matrices $W^{t}_{conv},W^{l}_{conv},W^{r}_{conv}$. For node $x_{i}$ in the window of a filter, its weight matrix is denoted as $W_{conv,i}$, which is a linear combination of those three weights above with corresponding coefficients $\eta_{i}^{t},\eta_{i}^{r}$, and $\eta_{i}^{l}$. Each layer’s output of TCN is calculated by the following formulas [30]:
$\eta_{i}^{t}=\frac{d_{i}-1}{d-1}\quad\eta_{i}^{r}$ $=(1-\eta_{i}^{t})\frac{p_{i}-1}{n-1},\quad\eta_{i}^{l}=(1-\eta_{i}^{t})(1-\eta_{i}^{r})$ (4)
$W_{conv,i}$ $=\eta_{i}^{t}W^{t}_{conv}+\eta_{i}^{l}W^{l}_{conv}+\eta_{i}^{r}W^{r}_{conv}$ (5)
$y$ $=tanh(\sum^{n}_{i=1}{W_{conv,i}\cdot x_{i}+b_{conv}})$ (6)
@@ IV-C Query Performance Prediction
CONCERTO’s GAT and TCN networks are trained together in a query-driven way since constructing the calibrated cost vector to the cost vector tree is differentiable. They are trained with the MSE loss function together with normalized ground-truth query latency as the label. To make the model converge more stably, all neural networks are connected by a residual connection path [42].
V Runtime Tracker
V-A Motivation
To develop a physical plan-level performance predictor for high-performance OLAP DBMSs, we need to collect detailed information to train and evaluate the predictor. However, such DBMSs’ parallel and dynamic execution mechanisms prevent us from using tools like ‘EXPLAIN’ to support query performance prediction for the following reasons.
Figure 3: Illustration of dynamic modification. The Join operator with a red background is modified from hash join to merge join and the MergeSort operators are added to the DAG in the probe execution phase.
To make an accurate query performance prediction, it is necessary to acquire detailed operators’ runtime information, such as their resource cost and runtime calling parameters. It would be hard to match each operator’s information with the pipeline structure given by the ‘EXPLAIN’ SQL command since most of them employ dynamic execution mechanisms that allow inserting and modifying operators including adding new execution paths and changing the algorithm used by operators such as sort and join, which is hard to foresee during performance prediction.
For the first type of action, assuming the DBMS finds that it mistakenly selected the merge join instead of the hash join algorithm, it will actively modify the join algorithm. Some DBMSs [20] achieve this by passing different types of algorithm parameters to the Join operator; the instance of the operator remains the same, but the operator will have different features and costs for most of the execution, which cannot be obtained by simply using the static pipeline. Other DBMSs, such as SparkSQL, change the algorithm by replacing the corresponding physical operator. In both situations, relying solely on the information provided in a static pipeline is unreliable.
For the second type of action, assuming that the DBMS decides to switch from internal sorting to external sorting, it will involve adding and replacing operators in the pipeline. Some DBMSs, like SparkSQL with the AQE mechanism, can even adjust the join order during execution. In such a situation, the execution paths of chunks can be completely changed.
Therefore, simply extracting static information from the planner is not enough; the runtime information, including calling parameters, pipeline modifications, and chunks execution information, is necessary for precise performance prediction. However, the pipeline exported by the ‘EXPLAIN’ statement not only does not contain any of those modification details, but it also cannot provide any runtime information about operators and chunks. Thus, it is difficult to use for QPP. Although it is impossible to obtain all pipeline modifications during the whole execution process while predicting query performance, inspired by ClickHouse’s probe phase [20], we developed a probe execution mode to launch, track, and collect the first several chunks’ execution information rapidly and obtain the pipeline structure, operators’ runtime features and pipeline modifications happened on the starting of the execution.
Considering the dynamic pipeline execution mechanism, those chunks can be generated by different execution paths that correspond to different features and costs. Thus, it would be a time-space graph with unaligned time steps. To collect that information, CONCERTO either embeds a variable number of chunks’ features from unaligned time slices into each operator of the pipeline or separates those chunks’ execution paths by constructing and processing a data flow tree. The former solution needs time-space graph convolution on a graph with many padded zero features, which could lead to resource waste and lower accuracy. Therefore, we chose the latter solution, which expands the execution path of multiple chunks involved in the probe execution in the spatial dimension and constructs a data-flow tree by combining them as the representation of the dynamic pipeline structure.
Besides, since CONCERTO is a multi-stage method that decouples the operator cost prediction, cost calibration, and query performance prediction, it also has to collect the training data of OCPs separately to avoid interference from other parallel operators.
With all those reasons above, we designed the Runtime Tracker to collect detailed runtime information to support more accurate QPP. For the convenience of discussion, we take ClickHouse, a representative high-performance OLAP DBMS, as an example to introduce the Runtime Tracker’s design in the following subsection. Note that such a design is portable between DBMSs since high-performance OLAP DBMSs have similar execution mechanisms such as shown in Table I, and we assume the target DBMS uses a DAG-based pipeline, which is a more general structure compared with the tree-shaped pipeline.
V-B Design of Runtime Tracker and Data Collection Process
Figure 4: Illustration of the structure of the Runtime Tracker. The Serial Executor in dotted lines needs to be modified from the DBMS’s parallel executor if the DBMS does not provide a serial execution mode like SparkSQL does.
As shown in Figure 4, the Runtime Tracker includes three parts:
  • Executor. The executor part can usually be regarded as a part of the DBMS’s execution engine. However, in order to obtain the OCPs’ training data without interference from the parallel execution of other operators, CONCERTO needs a serial executor. Some DBMSs like SparkSQL provide such a serial execution mode, but some like ClickHouse do not. For such DBMSs, it is necessary to create a serial executor from their parallel executor. Such modification is usually simple since most DBMSs with parallel executors provide a knob for the degree of parallelism (DOP).
  • Execution Mode. CONCERTO uses a full-collection mode to collect the training data for Stage 2 and Stage 3. Since CONCERTO collects the chunk-level information that cannot be fully obtained during inference, it employs a probe execution mode to obtain a part of the dynamically changing pipeline information.
  • Cost Logger. CONCERTO uses very detailed low-level information about the physical pipeline as features, including the operator’s cost, calling parameters, system resource utilization, and chunk execution information. The information is finally collected by Cost Logger, which involves some simple wrappers and modifications of the DBMS’s physical operators and data chunk class.
Data for OCPs. The features and execution costs of each single operator that is used to train the OCPs should be collected offline. Since the OCPs aim to predict the cost of each operator without interference from the parallel execution of other operators, their training data should be collected separately by a serial executor. To collect enough data with as few queries as possible, we run Runtime Tracker in full-collection mode, which collects every operator’s runtime features and costs on every data chunk. Although this will cause extra data collecting and training costs compared to using the history workload log, the overhead would not be too high because the OCPs are small MLP networks, and they only need a single-time data collection for a deployed DBMS.
We wrap the work() function of operators with our Cost Logger to record the operator’s calling parameters, system resource utilization, and resource cost during execution. The Cost Logger uses ioctl library to obtain the CPU instructions, CPU cycles, cache references, and cache misses. It uses the cpuinfo file to obtain the CPU frequency and uses proc/stat to obtain the CPU utilization. The Cost Logger also uses the proc/meminfo file of Linux to measure the memory cost and uses the /proc/PID/io file to measure the I/O read and write cost. After applying these modifications, the serial executor can be regarded as degenerating into a volcanic model. Then we execute workloads and collect each operator’s features and the costs of executing each data chunk. The data collection is efficient because there are dozens to hundreds of data chunks for each operator in one query.
Data for Cost Calibrator. As discussed above, CONCERTO solves the dynamic modification of pipeline and calling parameters during cost calibration and query performance prediction by expanding the runtime execution paths in the spatial dimension. This is achieved by tracking the execution path of chunks in the probe execution mode. During offline training data collection, CONCERTO will set the executor and the tracker to the probe execution mode and make a quick, short execution till the first few result chunks are collected. During this procedure, all intermediate chunks’ execution paths and operators’ calls are traced, and the potential dynamic modifications are also logged. The Graph Constructor will use this information in Stage 2 to construct a data-flow tree.
To track and log this information, we modified the code of the chunk and operators. We use the address of each operator instance as its ID to distinguish them. Once chunks are generated or processed, the corresponding operator will append its ID to the transform_addr list of the chunks. During the processing of chunks, Cost Logger would trace the costs and log them. The address of those cost records will also be added to the chunks’ record_addr list. By analyzing those records, CONCERTO can construct the data-flow tree for cost calibration and query performance prediction. The detailed construction algorithm will be introduced in  subsection IV-B.
Taking the TPC-DS workload as an example, the probe execution phase’s overhead is about 155ms. With this extra cost, we can obtain the data flow and partial dynamic added execution paths to support more accurate performance prediction. Since high-performance OLAP DBMSs are designed for processing complex OLAP queries that take hours to execute, the overhead is relatively small. Besides, this overhead is eliminable by marking those tuples that had been processed during probe execution and reading the result to skip them during the latter query execution.
This approach appears to have a limitation in its inability to capture all dynamic modifications occurring throughout the entirety of query execution, particularly those stemming from system load variations, since they are unpredictable. However, it remains effective in detecting modifications arising from estimation errors, such as cardinality errors. These types of adjustments typically occur during the early stages of query execution, which are adequately addressed by the probe phase.
VI Experiments
@@
In this section, we evaluate CONCERTO’s performance and compare it with other baseline methods. Since ClickHouse supports all the mechanisms as shown in Table I, we consider it one of the most representative high-performance OLAP DBMSs. Thus, we implement CONCERTO on ClickHouse for evaluation.
VI-A Experimental Settings
@@
Baselines. Since previous works already applied full comparison with conventional methods such as SVM and RBF [11, 34, 35] and proved learned QPP approaches outperformed conventional approaches on average accuracy, we will focus on the deep learning-based methods proposed in recent years. All baselines compared are listed as follows:
  1. 1.
@@
    QPPNet [11]: The QPPNet is an operator-level method that builds a cost model with a unified input and output interface for each operator and combines them according to the structure of the query plan tree to predict the latency. This method is considered to be a flexible model that can adapt to all different plan templates.
  2. 2.
@@
    QueryFormer [10]: The QueryFormer is a variant of the Transformer, which provides an attention mechanism on tree-shaped data and can capture a high-level expression of query plans to support downstream tasks such as QPP.
  3. 3.
@@
    RAAL [17]: The RAAL uses an attention-based resource-aware network to capture the resource allocation in SparkSQL and predict the query latency. The application scenario is quite similar to high-performance OLAP DBMSs’ parallel pipeline, so we include it as a competitive baseline. We implement it by following the formula and instructions in its paper.
  4. 4.
@@
    GCN  [18]: The GCN is a well-known network that is proposed to perform convolution on graph data. Since both DAG and tree-shaped pipelines can be regarded as a graph, we use GCN to perform end-to-end QPP as a baseline.
  5. 5.
@@
    TCN  [30]: The TCN is a common and widely used network in the domain related to query execution and query optimization [4], which introduces a convolution operation to the tree-shaped data. Thus, we included it as a baseline.
@@
We used their open-source code and modified it to adapt the data collected by CONCERTO’s Runtime Tracker from ClickHouse due to ClickHouse’s lack of a tool to export a detailed enough physical pipeline for baselines such as QPPNet and QueryFormer. In fact, CONCERTO’s Runtime Tracker provides even more comprehensive features than more DBMSs’ original implementation. For RAAL, since it is an approach designed for SparkSQL that considers the resource competition, which is also included in CONCERTO’s scenario, it can share the same training and testing data with CONCERTO (only in a different format). Thus, the comparison is carefully designed and implemented to ensure fairness.
@@
In addition to those baselines above, we also performed ablation experiments on CONCERTO to evaluate each component’s performance with the following setups:
  1. 1.
@@
    CONCERTO w/o ResAttn: This combination only removes the resource attention mechanism of CONCERTO and uses the expanded meta-competition matrix only as GAT’s input.
  2. 2.
@@
    CONCERTO w/o OCP: This combination removes the OCP module only and keeps the other modules to evaluate the impact of the OCP module on the QPP performance.
@@
Datasets. Followed by QPPNet [11] and MB2 [9], we compare the performance on TPC-H and TPC-DS benchmarks with $scale\_factor=1$ These are two standard benchmarks designed specifically for evaluating OLAP database performance and are widely used in the industry.
@@
Workloads. We use templates from TPC-H and TPC-DS for training and testing. The number of generated queries of each workload is about 10,000 (slightly fluctuating based on the number of query templates because we generate the same number of queries for each template). Among those queries, we evenly draw all the queries of 4 templates (templates 3, 7, 11, 6) from TPC-H (23.5%) and 9 templates (templates 3, 9, 15, 21, 27, 33, 39, 45, 51, 57) from TPC-DS (15.52%) by the index as test workloads that will not appear in the training workloads. Note that the testing workloads only use the data collected by the probe execution mode as input.
@@
Metric. We use the Q-Error as the metric to measure the accuracy of QPP, which is widely used in many previous QPP and cardinality estimation methods [11, 15, 36]. The definition of Q-Error is listed as follows:
Q-Error $=\frac{\max{(prediction,actual)}}{\min{(prediction,actual)}}$ (7)
@@
Model Architecture. Considering the inference speed, CONCERTO uses relatively shallow networks, such as a 2-layer GAT and a 3-layer TCN, and all OCPs are MLPs with 3 hidden layers.
@@
Hardware. All training and testing workloads are generated on AMD R9-7945HX, 32GB memory, and RTX4060 with a controllable environment that has no interference from other processes. We use a public server with 6 RTX A6000 (48GB memory) GPUs, 2 Intel(R) Xeon(R) Silver 4210R CPUs, and 504GB RAM to perform all evaluations on CONCERTO, baselines, and ablation experiments.
VI-B QPP Accuracy
TABLE II: Accuracy comparison of baselines, CONCERTO, and ablation studies.